An aeroplane takes off at an angle of 60 degree to the horizontal if the components of the velocity along the horizontal is 200 m/s what is the actual velocity? Also find the vertical components of its velocity


1
Expert's answer
2021-11-29T11:42:05-0500

"v_x=v\cdot\cos60°\to v=v_x/\cos60°=200/\cos60°="


"=400\ (m/s)"


"v_y=v\cdot\sin60°=400\cdot\sin60°=346\ (m/s)"